OpenStudy (anonymous):

I need some algebra help! 1. d^2 + 18d + 81 A) (d+9)(d-9) B) (d+9)^2 C) (d-81)(d-1) D) (d-9)^2 2. d^2 - 13d + 30 A) (d+10)(d+3) B) (d+10)(d-3) C) (d-10)(d+3) D) (d-10)(d-3) 3. r^2 - 49 A) (r-7)(r+7) B) (r+7(r+7) C) (r-7)(r-7) D) (r-7)(r+9)

OpenStudy (jango_in_dtown):

1. d^2 + 18d + 81=d^2 +9d +9d +81=d(d+9)+9(d+9)=(d+9)^2

OpenStudy (jango_in_dtown):

hints 81=9*9

OpenStudy (hwyl):

LOOKS LIKE YOU NEED TO FIGURE OUT THE MIDDLE AND LAST TERM THINK OF TWO NUMBERS THAT ADDS UP TO 18 AND YIELDS 81 WHEN MULTIPLIED TO EACH OTHER

OpenStudy (anonymous):

A C B C

OpenStudy (anonymous):

@jango_IN_DTOWN So number 1 is B?

OpenStudy (anonymous):

ignore the last c ;)

OpenStudy (jango_in_dtown):

yes

OpenStudy (jango_in_dtown):

2.d^2 - 13d + 30 i will give hints -13=-10+(-3) and 30=(-10)*(-3)

OpenStudy (jango_in_dtown):

3. 49=7^2 and so use the formula of a^2-b^2= (a+b)(a-b)